OpenStudy (yorel_1999):

Is the following sentence a statement or not a statement? He is six years old.

OpenStudy (ashely14):

i say statement because they are saying he is a six year old boy

likeabossssssss (likeabossssssss):

welcome to os

RhondaSommer (rhondasommer):

green or pink?

OpenStudy (yorel_1999):

Thank you Ashley. And thank you Likeaboss for welcoming me. I Just became a member minutes ago.

OpenStudy (ashely14):

green or pink what

OpenStudy (ashely14):

congrats @yorel_1999

RhondaSommer (rhondasommer):

yorel just answer. which one is better trust me it's about something good :D

OpenStudy (yorel_1999):

Pink?

OpenStudy (ashely14):

if you need more help just tag me @yorel_1999

RhondaSommer (rhondasommer):

\(\Huge\color{violet}{\mathbb{WELCOME}}\) \(\Huge\color{magenta}{\mathbb{TO}}\) \(\Huge\color{violet}{\mathbb{OPENSTUDY}}\)

OpenStudy (ashely14):

did you need help with anything else yorel

OpenStudy (yorel_1999):

Thank You Soooo Much Ashley. I Am Taking Online Classes For Algebra And Geometry Because I am Recovering. I am Not Very good With Math. :/ ANd Thank You Soooo Much Rhonda For The Great Big Welcome! ^_^ Pink Is My Favorite Color.

OpenStudy (ashely14):

np hope you recover soon

OpenStudy (yorel_1999):

A tautology is: always false always true sometimes true sometimes false